Calculate the proportion of F1 hybrid vigor retained at equilibrium for a four-way rotation using purebred A, B, C, and D sires.
Answer #1

Answer

%RHV = (m (2^n-1) -1) / (m (2^n-1)) x 100

n= number of sire types

m = number of breeds present in each sire type

%RHV = (2(2^4-1)-1)/ (2(2^4-1)) x 100

= (29 / 30) X 100

= 96.66%